OpenStudy (anonymous):

Solve for x. 3x - 1 = 20 -7 -19/3 19/3 7

OpenStudy (mew55):

add one to both sides 3x-1+1=20+1 3x=21 divide both sides by 3 3x/3=21/3 x=7

OpenStudy (anonymous):

thank you @mew55